OpenStudy (anonymous):

A bridge across a valley is 150 m in length. The valley walls make angles of 60° and 54° with the bridge that spans it, as shown. How deep is the valley, to the nearest metre?

OpenStudy (anonymous):

is it 72.88 aprox??

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):

well sorry !!! thats as good as i can do :( good luck !!

OpenStudy (anonymous):

oh lol thanks anyways :)

OpenStudy (anonymous):

this is the diagram

OpenStudy (lastdaywork):

Use x cot60 + x cot54 = 150

OpenStudy (lastdaywork):

where x is the depth of the valley..

random231 (random231):

yeah lastday is right

OpenStudy (anonymous):

Okay but what's "cot"

OpenStudy (anonymous):

... is it the same as "cosine"

OpenStudy (gudden):

no... cot = 1/tan

OpenStudy (lastdaywork):

cot = cos / sin As the question says - ".. to the nearest metre.." you can use cot 53 instead of cot 54

OpenStudy (anonymous):

oh okay... thank you guys so much :)

OpenStudy (gudden):

@LastDayWork now... my problem ... plz.... help me...out....

OpenStudy (anonymous):

The valley's depth is 115 meters. Refer to the attached solution using Mathematica.

OpenStudy (lastdaywork):

Okay;now I remember...there was a formula for cos 36 too

OpenStudy (anonymous):

@tania98 Thank you for the medal.